OpenStudy (anonymous):

The Bangor Coat Factory found that the number of coats sold N varies directly as their advertising budget A and inversely as the price of each coat P. The Bangor Coat Factory sold 10,800 coats when $30,000 was spent on advertising and the price of a coat was set at $50. Determine the number of coats sold when the amount spent on advertising is increased to $70,000 THATS THE QUESTION 3888 coats,24480,1400,25200 AND THOSE ARE THE MULTIPLE CHOICES

OpenStudy (gw2011):

The answer is 25,200. This can be solved by using proportion. Set up the proportion as follows: (10,800/30,000)=(x)/(70,000) By cross multiplying you get: (30,000)(x)=(70,000)(10,800) x=[(70,000)(10,800)]/30,000 x=25,200 coats